# Density Descent for Diversity Optimization

### v1

@@
Abstract
Diversity optimization seeks to discover a set of solutions that elicit diverse features. Prior work has proposed Novelty Search (NS), which, given a current set of solutions, seeks to expand the set by finding points in areas of low density in the feature space. However, to estimate density, NS relies on a heuristic that considers the $k$-nearest neighbors of the search point in the feature space, which yields a weaker stability guarantee. We propose Density Descent Search (DDS), an algorithm that explores the feature space via gradient descent on a continuous density estimate of the feature space that also provides stronger stability guarantee. We experiment with DDS and two density estimation methods: kernel density estimation (KDE) and continuous normalizing flow (CNF). On several standard diversity optimization benchmarks, DDS outperforms NS, the recently proposed MAP-Annealing algorithm, and other state-of-the-art baselines. Additionally, we prove that DDS with KDE provides stronger stability guarantees than NS, making it more suitable for adaptive optimizers. Furthermore, we prove that NS is a special case of DDS that descends a KDE of the feature space.
1 Introduction
@@
There is a range of applications where searching directly for behavioral diversity results in finding suboptimal solutions that act as “stepping stones,” mitigating convergence to local optima (Gaier et al.,, 2019). A classic example is the problem of training an agent to reach a target position in a deceptive maze (Lehman and Stanley,, 2011). There, directly minimizing the distance between the agent’s final position and a target goal causes the agent to get stuck. On the other hand, the problem can be solved by ignoring the objective of directly reaching the target and instead attempting to find a diverse range of agents, each of which reaches a different region of the maze.
@@
We refer to the problem of finding a range of solutions that are diverse with respect to prespecified features as diversity optimization (DO). We characterize DO as a special instance of quality diversity optimization (QD) (Pugh et al.,, 2016). Like DO, QD seeks a diverse set of solutions, but QD also considers an objective function that the solutions must optimize. For instance, in deceptive maze, one could add an energy consumption objective so that the goal is to find a set of agents that minimize energy consumption while moving to diverse regions in the maze. Thus, DO is an instance of QD where the objective value of all solutions is constant.
Figure 1: We propose density descent search (DDS) for solving diversity optimization (DO) problems. DDS first draws solutions from a Gaussian $\mathcal{N}({\bm{\mu}},{\bm{\Sigma}})$. After computing the solution features (in this case, the final position of the robot in a maze), DDS ranks solutions by density. This density ranking is passed to CMA-ES, which updates the search distribution to sample solutions with lower density on the next iteration. Concurrently, solutions are stored in a buffer that forms the basis for density estimates, and in a passive archive that tracks all discovered solutions.
@@
In DO and QD, discovering solutions with new features is challenging because the mapping from solutions to features is complex and non-invertible (Chenu et al.,, 2021). For example, in maze exploration, it is not known a priori how to produce an agent that travels to a given $(x,y)$ position.
@@
A classic method applied to DO is Novelty Search (NS) (Lehman and Stanley,, 2011). NS retains an archive of previously found individual solutions. It aims to expand the archive by finding solutions that are far in the feature space from existing solutions in the archive. Specifically, NS optimizes for solutions with a high novelty score, where novelty score is the average distance in feature space from a solution to its $k$-nearest neighbors in the archive.
@@
While the novelty score is framed in terms of distance, we study its interpretation as an approximation of density (Chenu et al.,, 2021), where density is directly proportional to the number of solutions in a region of feature space. A high novelty score indicates that a solution’s features are far from those of its $k$-nearest neighbors in the archive, i.e., it is located in an area of the feature space with low density.
@@
An alternative approach to DO is to apply general-purpose QD algorithms and assume that all solutions in the search space have identical quality, i.e., equal to some constant. Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) (Fontaine and Nikolaidis,, 2023) is a state-of-the-art QD algorithm that optimizes for archive improvement with the CMA-ES (Hansen,, 2016) optimizer. CMA-MAE retains a discrete archive of solutions in feature space. When applied to solve a DO problem by ignoring the solution quality, this archive naturally becomes a histogram that represents the distribution of solutions in feature space, with lower values of the histogram indicating lower density. CMA-MAE then performs density descent on this histogram, where it continually seeks to fill the areas of low density.
@@
Our key insight is to overcome the drawbacks of current density estimation methods in DO by introducing continuous, stable approximations of the solution density in feature space. This insight results in the following contributions: (1) We propose two new algorithms that query continuous density estimates to guide an optimizer, in our case CMA-ES, to search for solutions that are diverse in the feature space (Sec. 4). The first algorithm, DDS-KDE, leverages non-parametric Kernel Density Estimation (KDE) (Chen,, 2017). The second algorithm, DDS-CNF, leverages continuous normalizing flow (Lipman et al.,, 2023) to learn the underlying density function. (2) We show that, when combined with a ranking-based optimizer like CMA-ES, NS reduces to a special case of DDS-KDE (Sec. 5). (3) We prove that KDE provides stronger algorithmic stability guarantees than novelty score (Sec. 5). (4) We demonstrate that our DDS algorithms outperform prior work on 3 out of 4 domains (Sec. 6). (5) We show that our algorithms perform well on multi-dimensional feature spaces, which currently present significant challenges to QD and DO algorithms.
2 Problem Definition
@@ 3.1 QD and DO Algorithms
Novelty Search (NS). The key insight of NS (Lehman and Stanley,, 2011) is to discover diverse solutions in feature space by optimizing for solutions that are “novel” with respect to a current set of solutions. Given a set of features $\mathcal{X}\subseteq\mathbb{R}^{m}$, the novelty of a solution ${\bm{\theta}}\in\mathbb{R}^{n}$ and its features ${\bm{y}}={\bm{\phi}}({\bm{\theta}})$ relative to $\mathcal{X}$ is encapsulated in its novelty score, denoted $\rho({\bm{y}};\mathcal{X})$:
$\rho({\bm{y}};\mathcal{X})=\frac{1}{k}\sum_{{\bm{y}}^{\prime}\in N_{k}({\bm{y}};\mathcal{X})}\mathrm{dist}({\bm{y}},{\bm{y}}^{\prime})$ (1)
@@
Multi-dimensional Archive of Phenotypic Elites (MAP-Elites). While NS was developed for DO, MAP-Elites (Mouret and Clune,, 2015) was developed for the general QD setting. Compared to the unstructured archive of NS, MAP-Elites divides the feature space into a predefined tessellation $T:\mathbb{R}^{m}\to\{1,\dots,l\}$, where $e\in\{1,\dots,l\}$ is a cell in the tessellation and $l$ is the total number of cells. Given a cell $e$, the MAP-Elites archive associates a solution ${\bm{\theta}}$ with cell $e$ if and only if the solution’s features are contained in $e$, i.e., $T({\bm{\phi}}({\bm{\theta}}))=e$. Moreover, MAP-Elites stores at most one solution for every cell in the tessellation.
@@
During a QD search, MAP-Elites first draws solutions from a predefined distribution, e.g., a Gaussian, and inserts the solutions into the archive. Subsequently, MAP-Elites generates and inserts new solutions by mutating existing archive solutions with a genetic operator, such as the Iso+LineDD operator (Vassiliades and Mouret,, 2018). Importantly, when solutions inserted into the archive land in the same tessellation cell, MAP-Elites only retains the solution with the greatest objective value. Thus, MAP-Elites gradually collects high-performing solutions that have diverse features. We apply MAP-Elites to DO by setting a constant objective $f(\cdot)=C$ over the solution space.
@@
The choice of tessellation in MAP-Elites can significantly impact scalability. The most common tessellation is a grid tessellation (Mouret and Clune,, 2015), which divides the feature space into equally-sized hyperrectangles. In high-dimensional feature spaces, grid tessellations require exponentially more memory due to the curse of dimensionality. Thus, a common alternative is the centroidal Voronoi tessellation (CVT) (Vassiliades et al.,, 2018), which divides the feature space into $l$ evenly-sized polytopes.
@@
Covariance Matrix Adaptation Evolution Strategy (CMA-ES). One recent line of QD algorithms has combined CMA-ES (Hansen,, 2016) with MAP-Elites. CMA-ES is a state-of-the-art single-objective optimizer that represents a population of solutions with a multivariate Gaussian $\mathcal{N}({\bm{\mu}},{\bm{\Sigma}})$. Each iteration, CMA-ES draws $\lambda$ solutions from the Gaussian. Based on the rankings (rather than the raw objectives) of the solutions, CMA-ES adapts the covariance matrix ${\bm{\Sigma}}$ to regions of higher-performing solutions. While CMA-ES is a derivative-free optimizer, it has been shown to approximate a natural gradient (Akimoto et al.,, 2010).
@@
Covariance Matrix Adaptation MAP-Elites (CMA-ME). The first work to integrate CMA-ES with MAP-Elites was CMA-ME (Fontaine et al.,, 2020). The key idea of CMA-ME is to optimize for archive improvement with CMA-ES. Namely, in addition to a MAP-Elites-style archive, CMA-ME maintains an instance of CMA-ES. The CMA-ES instance samples solutions from a Gaussian, and the solutions are ranked based on how much they improve the archive, e.g., solutions that found new cells in the archive are ranked high, while those that were not added at all are ranked low. With this improvement ranking, CMA-ES adapts the Gaussian to sample solutions that will further improve the archive. Note that in DO settings, since the objective is a constant value, the improvement ranking places solutions that filled a new cell ahead of solutions that did not add a new cell.
@@
Covariance Matrix Adaptation MAP-Annealing (CMA-MAE). One limitation of CMA-ME is that it focuses too much on exploring for new cells in feature space rather than optimizing the objective (Tjanaka et al.,, 2022). CMA-MAE (Fontaine and Nikolaidis,, 2023) addresses this problem by introducing an archive learning rate $\alpha$. When $\alpha=1$, CMA-MAE maintains the same exploration behavior as CMA-ME, but when $\alpha=0$, CMA-MAE focuses solely on optimizing the objective, like CMA-ES. For $10<\alpha<$ 1, CMA-MAE blends between these two extremes, enabling it to both explore feature space and optimize the objective.
@@
In DO settings, where the objective is constant ($f(\cdot)=C$), CMA-MAE naturally performs density descent (Fontaine and Nikolaidis,, 2023). Intuitively, the archive becomes a histogram that represents how many solutions have been found in each region of feature space. Lower values of the histogram indicate lower density of solutions, i.e., if a histogram bin has a low value, few solutions have been discovered in that region of feature space. CMA-MAE then seeks to descend the histogram by searching for solutions that fill the low-valued histogram bins.
3.2 Density Estimation
@@
Kernel Density Estimation (KDE). KDE is a non-parametric density estimation method that does not make any assumptions on the underlying probability density distribution from which samples are drawn (Chen,, 2017). Given a set of features $\mathcal{X}\subseteq\mathbb{R}^{m}$, bandwidth parameter $h$, and kernel function $K(\cdot)$, KDE computes the density function $\hat{D}:\mathbb{R}^{m}\to\mathbb{R}$ for a given feature ${\bm{y}}\in\mathbb{R}^{m}$ as:
$\hat{D}_{h}({\bm{y}};\mathcal{X})=\frac{1}{|\mathcal{X}|h}\sum_{{\bm{y}}^{\prime}\in\mathcal{X}}K\left(\frac{\lVert{\bm{y}}-{\bm{y}}^{\prime}\rVert}{h}\right)$ (2)
Prior work has thoroughly studied the problem of selecting bandwidth that accurately estimate the underlying density function (Silverman,, 1986; Rudemo,, 1982; Bowman,, 1984). While accurate density estimation is important for KDE, we show in Theorem 5.2 that larger bandwidth results in more stable density estimates (albeit at the cost of accuracy), which is beneficial to the proposed algorithms in Sec. 4.
When optimizing a density estimate (e.g., with gradient descent), KDEs have several advantages over histograms. First, the shape of a histogram depends on its bin size (Weglarczyk,, 2018). Second, the binning procedure leads to a discontinuous optimization landscape and flat gradient signals inside each bin (Weglarczyk,, 2018). In contrast, KDEs produce a smooth, continuous density function based on the location of the samples in the underlying distribution (Chen,, 2017).
Continuous Normalizing Flow (CNF). CNF (Lipman et al.,, 2023) is a generative modeling method that constructs a diffusion path between a simple distribution (e.g., a Gaussian distribution) and an unknown distribution. The diffusion path describes a mapping from a point on the simple distribution to a corresponding point on the unknown distribution such that their probability densities are roughly equal. CNFs enable sampling from probability distributions where direct sampling is difficult (e.g., distributions of images). This is accomplished by sampling from the simple distribution and transforming the sample via the learned diffusion path of the CNF. However, we utilize CNF to estimate the density of feature space for our proposed algorithms in Sec. 4.
4 Density Descent Search
We present Density Descent Search (DDS), a DO algorithm that efficiently explores the feature space by leveraging continuous density estimates. DDS maintains a buffer of features ${\mathcal{B}}$ that represents the distribution of features discovered so far. Based on this buffer, DDS models a density estimation of the feature distribution, and by querying this density estimate, DDS guides an adaptive optimizer to discover solutions in less-dense areas of feature space. As it searches for such solutions, DDS expands the discovered set of features. Below, we describe the components of DDS.

### v2

@@
Abstract.
Diversity optimization seeks to discover a set of solutions that elicit diverse features. Prior work has proposed Novelty Search (NS), which, given a current set of solutions, seeks to expand the set by finding points in areas of low density in the feature space. However, to estimate density, NS relies on a heuristic that considers the $k$-nearest neighbors of the search point in the feature space, which yields a weaker stability guarantee. We propose Density Descent Search (DDS), an algorithm that explores the feature space via CMA-ES on a continuous density estimate of the feature space that also provides a stronger stability guarantee. We experiment with DDS and two density estimation methods: kernel density estimation (KDE) and continuous normalizing flow (CNF). On several standard diversity optimization benchmarks, DDS outperforms NS, the recently proposed MAP-Annealing algorithm, and other state-of-the-art baselines. Additionally, we prove that DDS with KDE provides stronger stability guarantees than NS, making it more suitable for adaptive optimizers. Furthermore, we prove that NS is a special case of DDS that descends a KDE of the feature space.
quality diversity, kernel density estimation, novelty search
journalyear: 2024copyright: rightsretainedconference: Genetic and Evolutionary Computation Conference; July 14–18, 2024; Melbourne, VIC, Australiabooktitle: Genetic and Evolutionary Computation Conference (GECCO ’24), July 14–18, 2024, Melbourne, VIC, Australiadoi: 10.1145/3638529.3654001isbn: 979-8-4007-0494-9/24/07ccs: Computing methodologies Search methodologiesccs: Mathematics of computing Density estimation
1. Introduction
Figure 1. We propose density descent search (DDS) for solving diversity optimization (DO) problems. DDS first draws solutions from a Gaussian $\mathcal{N}({\bm{\mu}},{\bm{\Sigma}})$. After computing the solution features (in this case, the final position of the robot in a maze), DDS ranks solutions by density. This density ranking is passed to CMA-ES, which updates the search distribution to sample solutions with lower density on the next iteration. Concurrently, solutions are stored in a buffer that forms the basis for density estimates, and in a passive archive that tracks all discovered solutions.
@@ 1. Introduction
There is a range of applications where searching directly for behavioral diversity results in finding suboptimal solutions that act as “stepping stones,” mitigating convergence to local optima (Gaier et al., 2019). A classic example is the problem of training an agent to reach a target position in a deceptive maze (Lehman and Stanley, 2011). There, directly minimizing the distance between the agent’s final position and a target goal causes the agent to get stuck. On the other hand, the problem can be solved by ignoring the objective of directly reaching the target and instead attempting to find a diverse range of agents, each of which reaches a different region of the maze.
@@
We refer to the problem of finding a range of solutions that are diverse with respect to prespecified features as diversity optimization (DO). We characterize DO as a special instance of quality diversity optimization (QD) (Pugh et al., 2016). Like DO, QD seeks a diverse set of solutions, but QD also considers an objective function that the solutions must optimize. For instance, in deceptive maze, one could add an energy consumption objective so that the goal is to find a set of agents that minimize energy consumption while moving to diverse regions in the maze. Thus, DO is an instance of QD where the objective value of all solutions is constant.
@@
In DO and QD, discovering solutions with new features is challenging because the mapping from solutions to features is complex and non-invertible (Chenu et al., 2021). For example, in maze exploration, it is not known a priori how to produce an agent that travels to a given $(x,y)$ position.
@@
A classic method applied to DO is Novelty Search (NS) (Lehman and Stanley, 2011). NS retains an archive of previously found individual solutions. It aims to expand the archive by finding solutions that are far in the feature space from existing solutions in the archive. Specifically, NS optimizes for solutions with a high novelty score, where novelty score is the average distance in feature space from a solution to its $k$-nearest neighbors in the archive.
@@
While the novelty score is framed in terms of distance, we study its interpretation as an approximation of density (Chenu et al., 2021), where density is directly proportional to the number of solutions in a region of feature space. A high novelty score indicates that a solution’s features are far from those of its $k$-nearest neighbors in the archive, i.e., it is located in an area of the feature space with low density.
@@
An alternative approach to DO is to apply general-purpose QD algorithms and assume that all solutions in the search space have identical quality, i.e., equal to some constant. Covariance Matrix Adaptation MAP-Annealing (CMA-MAE) (Fontaine and Nikolaidis, 2023) is a state-of-the-art QD algorithm that optimizes for archive improvement with the CMA-ES (Hansen, 2016) optimizer. CMA-MAE retains a discrete archive of solutions in feature space. When applied to solve a DO problem by ignoring the solution quality, this archive naturally becomes a histogram that represents the distribution of solutions in feature space, with lower values of the histogram indicating lower density. CMA-MAE then performs density descent on this histogram, where it continually seeks to fill the areas of low density.
@@
Our key insight is to overcome the drawbacks of current density estimation methods in DO by introducing continuous, stable approximations of the solution density in feature space. This insight results in the following contributions: (1) We propose density descent search (DDS; Fig. 1), an algorithm that queries continuous density estimates to guide an optimizer, in our case CMA-ES, to search for solutions that are diverse in the feature space (Sec. 4). We propose two variants of DDS: DDS-KDE leverages non-parametric Kernel Density Estimation (KDE) (Chen, 2017), while DDS-CNF learns the underlying density function with continuous normalizing flow (Lipman et al., 2023). (2) We show that, when combined with a ranking-based optimizer like CMA-ES, NS reduces to a special case of DDS-KDE (Sec. 5). (3) We prove that KDE provides stronger algorithmic stability guarantees than novelty score (Sec. 5). (4) We demonstrate that our DDS algorithms outperform prior work on 3 out of 4 domains (Sec. 6). (5) We show that our algorithms perform well on multi-dimensional feature spaces, which currently present significant challenges to QD and DO algorithms.
2. Problem Definition
@@ 3.1. QD and DO Algorithms
Novelty Search (NS). The key insight of NS (Lehman and Stanley, 2011) is to discover diverse solutions in feature space by optimizing for solutions that are “novel” with respect to a current set of solutions. Given a set of features $\mathcal{X}\subseteq\mathbb{R}^{m}$, the novelty of a solution ${\bm{\theta}}\in\mathbb{R}^{n}$ and its features ${\bm{y}}={\bm{\phi}}({\bm{\theta}})$ relative to $\mathcal{X}$ is encapsulated in its novelty score, denoted $\rho({\bm{y}};\mathcal{X})$:
(1) $\rho({\bm{y}};\mathcal{X})=\frac{1}{k}\sum_{{\bm{y}}^{\prime}\in N_{k}({\bm{y}};\mathcal{X})}\mathrm{dist}({\bm{y}},{\bm{y}}^{\prime})$
@@
While we consider general optimization domains in this work, we note that prior works (Conti et al., 2018) have specialized NS for reinforcement learning domains, particularly the case when no feature function is assumed (Paolo et al., 2020; Cully, 2019).
@@
Multi-dimensional Archive of Phenotypic Elites (MAP-Elites). While NS was developed for DO, MAP-Elites (Mouret and Clune, 2015) was developed for the general QD setting. Compared to the unstructured archive of NS, MAP-Elites divides the feature space into a predefined tessellation $T:\mathbb{R}^{m}\to\{1,\dots,l\}$, where $e\in\{1,\dots,l\}$ is a cell in the tessellation and $l$ is the total number of cells. Given a cell $e$, the MAP-Elites archive associates a solution ${\bm{\theta}}$ with cell $e$ if and only if the solution’s features are contained in $e$, i.e., $T({\bm{\phi}}({\bm{\theta}}))=e$. Moreover, MAP-Elites stores at most one solution for every cell in the tessellation.
@@
During a QD search, MAP-Elites first draws solutions from a predefined distribution, e.g., a Gaussian, and inserts the solutions into the archive. Subsequently, MAP-Elites generates and inserts new solutions by mutating existing archive solutions with a genetic operator, such as the Iso+LineDD operator (Vassiliades and Mouret, 2018). Importantly, when solutions inserted into the archive land in the same tessellation cell, MAP-Elites only retains the solution with the greatest objective value. Thus, MAP-Elites gradually collects high-performing solutions that have diverse features. We apply MAP-Elites to DO by setting a constant objective $f(\cdot)=C$ over the solution space.
@@
The choice of tessellation in MAP-Elites can significantly impact scalability. The most common tessellation is a grid tessellation (Mouret and Clune, 2015), which divides the feature space into equally-sized hyperrectangles. In high-dimensional feature spaces, grid tessellations require exponentially more memory due to the curse of dimensionality. Thus, a common alternative is the centroidal Voronoi tessellation (CVT) (Vassiliades et al., 2018), which divides the feature space into $l$ evenly-sized polytopes.
@@
Covariance Matrix Adaptation Evolution Strategy (CMA-ES). One recent line of QD algorithms has combined CMA-ES (Hansen, 2016) with MAP-Elites. CMA-ES is a state-of-the-art single-objective optimizer that represents a population of solutions with a multivariate Gaussian $\mathcal{N}({\bm{\mu}},{\bm{\Sigma}})$. Each iteration, CMA-ES draws $\lambda$ solutions from the Gaussian. Based on the rankings (rather than the raw objectives) of the solutions, CMA-ES adapts the covariance matrix ${\bm{\Sigma}}$ to regions of higher-performing solutions. While CMA-ES is a derivative-free optimizer, it has been shown to approximate a natural gradient (Akimoto et al., 2010).
@@
Covariance Matrix Adaptation MAP-Elites (CMA-ME). The first work to integrate CMA-ES with MAP-Elites was CMA-ME (Fontaine et al., 2020). The key idea of CMA-ME is to optimize for archive improvement with CMA-ES. Namely, in addition to a MAP-Elites-style archive, CMA-ME maintains an instance of CMA-ES. The CMA-ES instance samples solutions from a Gaussian, and the solutions are ranked based on how much they improve the archive, e.g., solutions that found new cells in the archive are ranked high, while those that were not added at all are ranked low. With this improvement ranking, CMA-ES adapts the Gaussian to sample solutions that will further improve the archive. Note that in DO settings, since the objective is a constant value, the improvement ranking places solutions that filled a new cell ahead of solutions that did not add a new cell.
@@
Covariance Matrix Adaptation MAP-Annealing (CMA-MAE). One limitation of CMA-ME is that it focuses too much on exploring for new cells in feature space rather than optimizing the objective (Tjanaka et al., 2022). CMA-MAE (Fontaine and Nikolaidis, 2023) addresses this problem by introducing an archive learning rate $\alpha$. When $\alpha=1$, CMA-MAE maintains the same exploration behavior as CMA-ME, but when $\alpha=0$, CMA-MAE focuses solely on optimizing the objective, like CMA-ES. For $10<\alpha<$ 1, CMA-MAE blends between these two extremes, enabling it to both explore feature space and optimize the objective.
@@
In DO settings, where the objective is constant ($f(\cdot)=C$), CMA-MAE naturally performs density descent (Fontaine and Nikolaidis, 2023). Intuitively, the archive becomes a histogram that represents how many solutions have been found in each region of feature space. Lower values of the histogram indicate lower density of solutions, i.e., if a histogram bin has a low value, few solutions have been discovered in that region of feature space. CMA-MAE then seeks to descend the histogram by searching for solutions that fill the low-valued histogram bins.
3.2. Density Estimation
@@
Kernel Density Estimation (KDE). KDE is a non-parametric density estimation method that does not make any assumptions on the underlying probability density distribution from which samples are drawn (Chen, 2017). Given a set of features $\mathcal{X}\subseteq\mathbb{R}^{m}$, bandwidth parameter $h$, and kernel function $K(\cdot)$, KDE computes the density function $\hat{D}:\mathbb{R}^{m}\to\mathbb{R}$ for a given feature ${\bm{y}}\in\mathbb{R}^{m}$ as:
(2) $\hat{D}_{h}({\bm{y}};\mathcal{X})=\frac{1}{|\mathcal{X}|h}\sum_{{\bm{y}}^{\prime}\in\mathcal{X}}K\left(\frac{\lVert{\bm{y}}-{\bm{y}}^{\prime}\rVert}{h}\right)$
Prior work has thoroughly studied the problem of selecting bandwidth that accurately estimate the underlying density function (Silverman, 1986; Rudemo, 1982; Bowman, 1984). While accurate density estimation is important for KDE, we show in Theorem 5.2 that larger bandwidth results in more stable density estimates (albeit at the cost of accuracy), which is beneficial to the proposed algorithms in Sec. 4.
When optimizing a density estimate (e.g., with gradient descent), KDEs have several advantages over histograms. First, the shape of a histogram depends on its bin size (Weglarczyk, 2018). Second, the binning procedure leads to a discontinuous optimization landscape and flat gradient signals inside each bin (Weglarczyk, 2018). In contrast, KDEs produce a smooth, continuous density function based on the location of the samples in the underlying distribution (Chen, 2017).
Continuous Normalizing Flow (CNF). CNF (Lipman et al., 2023) is a generative modeling method that constructs a diffusion path between a simple distribution (e.g., a Gaussian distribution) and an unknown distribution. The diffusion path describes a mapping from a point on the simple distribution to a corresponding point on the unknown distribution such that their probability densities are roughly equal. CNFs enable sampling from probability distributions where direct sampling is difficult (e.g., distributions of images). This is accomplished by sampling from the simple distribution and transforming the sample via the learned diffusion path of the CNF. However, we utilize CNF to estimate the density of feature space for our proposed algorithms in Sec. 4.
4. Density Descent Search
We present Density Descent Search (DDS), a DO algorithm that efficiently explores the feature space by leveraging continuous density estimates. DDS maintains a buffer of features ${\mathcal{B}}$ that represents the distribution of features discovered so far. Based on this buffer, DDS models a density estimation of the feature distribution, and by querying this density estimate, DDS guides an adaptive optimizer to discover solutions in less-dense areas of feature space. As it searches for such solutions, DDS expands the discovered set of features. We provide an overview of DDS in Fig. 1 and describe the components of DDS below.